OpenStudy (anonymous):

Two legs of a triangle measure 16 m and 34 m. How many meters long is the hypotenuse? Round your answer to the nearest tenth.

OpenStudy (anonymous):

you can use the pythagorean theorem a^2 + b^2 = c^2 (length of one side ) squared + (length of one side ) squared = hypotenuse

Directrix (directrix):

(length of one side ) squared + (length of one side ) squared = (hypotenuse)^2 16^2 + 34^2 = h^2